# V-VAE: A Variational Auto Encoding Framework Towards Fine-Grained Control over Human-Like Chat

### v1

@@ A.1 Derivation of the Objective
Existing dialogue datasets suffer from several limitations. Many lack fine-grained personal information  (Zhang et al., 2018a); others, often scraped from the web, contain noisy content and highly diverse topics (Wu et al., 2016). Some datasets lack long-term multi-turn interaction, exhibit short conversation depth (Li et al., 2017), or are restricted to narrow topical domains  (Zhang et al., 2018b). More recently, with the emergence of large language models (LLMs), many datasets have been constructed via instruction-following generation (e.g., GPT-based synthetic data) to suit specific downstream tasks (Su et al., 2020). However, such data often lacks authentic human attributes. In light of the absence of publicly available datasets that offer both multi-turn human-chatbot interaction and rich persona-related signals, we construct a new dataset to address these gaps.
A.3 Case Study

### v2

@@ 5 Conclusion
Acknowledgment
This work was supported by the National Key R&D Program of China (No. 2023YFC3305600).
@@ A.2 Human–Chatbot Chat Corpus
Existing dialogue datasets suffer from several limitations. Many lack fine-grained personal information  (zhang2018personalizing); others, often scraped from the web, contain noisy content and highly diverse topics (wu2016sequential). Some datasets lack long-term multi-turn interaction, exhibit short conversation depth (li-etal-2017-dailydialog), or are restricted to narrow topical domains  (zhang2018dua). More recently, with the emergence of large language models (LLMs), many datasets have been constructed via instruction-following generation (e.g., GPT-based synthetic data) to suit specific downstream tasks (su2020moviechats). However, such data often lacks authentic human attributes. In light of the absence of publicly available datasets that offer both multi-turn human-chatbot interaction and rich persona-related signals, we construct a new dataset to address these gaps.
A.3 Annotators’ Demographic Background
Our annotation cohort comprised 30 volunteers recruited from top-tier universities in Beijing (mean age=22.3±1.7 years), all native Mandarin speakers aligned with the linguistic context of the Chinese social platform used for human-Agent interactions. While this homogeneous demographic was intentionally selected to control cultural variables during the initial validation of fine-grained conversational patterns—consistent with common practices in foundational studies focusing on core user cohorts—we recognize its limitations in assessing broader sociocultural generalizability. To mitigate potential bias, we ensured geographic diversity in annotators’ regional origins (hometowns spanning 16 major Chinese cities), deliberately counterbalancing Beijing-centric educational enrollment with nationwide demographic representation.
CP EC HM ED KRG OD DS IC RC SF
DS-7B-base FT 1.85 4.4 0.4 0.5 12.9 30.2 3.4 46.9 39.0 21.4 24.3
P+FT 1.75 45.8 30.0 4.9 21.9 43.8 0.7 49.6 39.9 26.3 27.5
SP+FT 1.77 11.3 2.9 1.6 17.7 40.8 0.0 47.6 36.2 23.2 22.2
DS-7B-chat FT 1.85 4.3 0.5 1.2 44.4 71.0 50.6 67.1 59.0 61.2 73.7
P+FT 1.75 46.2 31.4 4.8 45.8 68.4 58.5 67.8 58.8 59.8 73.5
SP+FT 1.77 10.3 3.6 1.1 46.8 69.1 54.9 65.9 59.5 60.6 71.9
chatglm3-6B FT 2.28 9.1 1.2 0.7 35.0 47.6 52.7 63.3 62.6 60.4 60.2
P+FT 2.16 37.0 22.0 4.9 31.1 35.7 52.7 60.3 58.2 50.2 55.4
SP+FT 2.18 10.6 2.7 1.2 32.0 37.9 54.2 60.7 58.5 53.9 58.4
Target (ref) - 8.6 6.4 2.2
A.4 Additional Experiment
As is shown in Table 6, we evaluated more models with three training methods on HumanChatBench and DialogBench(ou2023dialogbench).The conclusions remain consistent with those reported in the main text.
A.5 Case Study